OpenStudy (anonymous):

If integral of f'(x) (f(x) dx = 6 and a>0 then [ f(a)^2)-(f(0))2]=?

OpenStudy (anonymous):

\[\int\limits_{0}^{a} f'(x) f(x) dx =6 \] and a>0 then [ f(a)^2)-(f(0))2]=?

OpenStudy (anonymous):

damn latex is not working for me

OpenStudy (anonymous):

\[\int_0^a udu=\frac{1}{2}\left(u^2(a)-u^2(0)\right)\]

OpenStudy (anonymous):

since this is equal to 6, your answer is 12

OpenStudy (anonymous):

no it's right , sorry I did a mistake. it supposed to be [(f(a))^2 - (f(0))^2)]

OpenStudy (anonymous):

thanks mate , I just wanted the idea how to solve it :') I appreciate it:D I have an exam tomorrow wish me the best, thanks again:D